# Risk-Aware Distributional Intervention Policies for Language Models

### v1

@@ 3. Headwise Interventions with Probabilistic Guarantees
where $\Delta_{\ell h}(a_{\ell h})=G^{\star}_{\ell h}a_{\ell h}+g^{\star}_{\ell h}$. In other words, each new headwise activation $\hat{a}_{\ell h}$ is computed based on three terms: the original headwise activations $a_{\ell h}$, the headwise intervention $\Delta_{h}(a_{\ell h})$, and the indicator value identifying if head $h$ and layer $\ell$ is predicted desirable or undesirable.
4. Experiments
@@
In this section, we present empirical evidence for the effectiveness of our method RADIANT. We evaluate RADIANT on the TruthfulQA benchmark [25], consisting of two tasks: the main task is the generation, and the secondary task is multiple choice. The generation task requires the model to generate an entire answer for each question using greedy autoregressive decoding. The accuracy and helpfulness of the answer are best assessed by humans. However, in almost all recent works in the field, including [24] and [50], this criterion is measured by an alternative large language model finetuned on the target dataset. The multiple-choice task contains candidate answers to each question, requiring the model to give probabilities. Higher probabilities for truthful answers yield higher scores.
4.1. Experimental Settings
Datasets. We evaluate and compare our method with other baselines using the TruthfulQA benchmark [25]. The TruthfulQA dataset is a Question-Answer dataset containing 817 questions that likely elicit false answers from humans due to common misconceptions. We follow the same data-processing used in [24] and [50] that splits the dataset into train/validation/test with the rate of $326/82/407$ questions and utilize two-fold cross-validation. Each question has an average length of nine words and has two sets of desirable and undesirable answers. Following [24], we separate the original dataset into 5918 question-answer pairs; each has a binary label, indicating desirability. Only pairs associated with questions in the training dataset are used to create our intervention policy, while those in the validation test are set aside for parameter tuning.
@@ 4.1. Experimental Settings
In addition, we also show the generalization of our method by conducting a transferability experiment on two other out-of-distribution datasets, including NQOpen [22] and TriviaQA [21]. Due to space constraints, the results are relegated to Appendix A.2.
Models. We implement our methods on various open-source pretrained Llama base models: Llama-7B [42], Llama2-chat-13B [43], and Llama3-8B [11]. Our method could be integrated with other methods as a tail component to efficiently elicit truthful answers from LMs. Therefore, we also used models fine-tuned for specific tasks to show the effectiveness of our approach.
@@
Hyperparameters. There are two pivotal hyperparameters in the RADIANT framework, namely $\alpha$ in probe loss (2), and $\Gamma=\Phi^{-1}(1-\gamma)$ in the computation of the intervention map (4). The discussion about their impact on RADIANT and how to select them is in Appendix A.1.
Baselines. We include baselines relevant to increasing truthfulness, listed as follows.
  • Supervised fine-tuning (SFT) is introduced in [31] as a component in an attempt to make LLMs align with human preference. Given a prompt, SFT encourages the model to generate desirable answers and reduce the likelihood of generating undesirable answers by optimizing cross-entropy loss. However, SFT’s requirement to finetune all LLM parameters demands substantial GPU resources for backpropagation operation. Due to computational constraints, we can only perform SFT on the GPT2-large, the smallest model in our experiments. The results are available in Table 9. This again highlights the advantages of inference time methods like ours: by avoiding gradient computation or backpropagation, they offer a lightweight, fast, versatile, and economical way to improve the performance of LLMs. This is especially useful in low-resource scenarios. Because Llama-7B is used as a base model for many of our experiments, we also include the results SFT on Llama-7B for comparison, but it is worth noting that this result is referred from the ITI paper [24]. Since our evaluation framework differs from ITI in terms of the GPT-judge and GPT-info models, which is attributed to the fact that these models in the ITI paper are no longer available in the OpenAI, the results may not be fair for comparison. From Table A.3, SFT achieves the best performance in terms of MC metrics and reaches a high score of True * Info and True. Regarding the True score, RADIANT still outperforms SFT in the individual and integrating versions with FSP, offering 38.73% and 40.41% correct answers, respectively. When combined with FSP, RADIANT achieves 35.36 % in True * Info score, surpassing SFT but requiring much less resources. For the implementation of SFT, we use the SFTTrainer framework from Hugging Face, one of the most popular frameworks for this algorithm. While we remained almost the default parameters proposed by the library, we had to tune many important parameters like learning rate, parameters of Adam optimizer, weight decay, and so on to get a consistent and stable fine-tuned model. Some important parameters for SFT are reported in the table below, while its best performance is represented in Table 9. This observation strongly supports the practicability of RADIANT, which only necessitates tuning two key hyper-parameters $\alpha$ in the probe loss (2), and $\Gamma=\Phi^{-1}(1-\gamma)$ in the computation of the intervention map (4). A thorough analysis of these parameters in an attempt to offer insights into their impact is presented in Appendix A.1. This section offers useful insights and detailed guidelines to select values for any new models. Furthermore, compared to other methods in the field, like ITI, we declared that the grid search on two hyper-parameters like ours is efficient and reasonable, so it is not harder to tune the hyper-parameters of RADIANT than other previous works.
    Parameter Value
    learning_rate 0.00002
    weight_decay 0
    adam_beta1 0.8
    adam_beta2 0.999
    adam_epsilon $1\times 10^{-8}$
    max_grad_norm 1
    batch_size 32
    epochs_num 5
    lr_scheduler_type linear
    Table 1. Parameter values for SFT.
@@
  • Inference-time Intervention (ITI [24]), the state-of-the-art method for finetuning-free intervention. The hyperparameters of the baseline follow their original paper [24] and their GitHub repository. [^1] https://github.com/likenneth/honest_llama/tree/master
  • Few-shot prompting (FSP) introduced in [2] showcases the effectiveness of 50-shot prompting in benchmark TruthfulQA.
@@
  • Instruction Fine-Tuning (IFT) [47, 9] is a popular fine-tuning approach to boost the truthfulness of language models. Two notable pretrained models in this direction, namely Alpaca-7B [39] and Vicuna-7B [8], are adopted for comparison.
@@
  • Representation Intervention Fine-tuning (RIFT) methods aim to adjust language model activations for improved truthfulness. However, they add extra parameters and require extensive computational resources for fine-tuning. We consider LOFiT [50] for comparison.
@@
  • Non-Linear Inference Time Intervention (NL-ITI) [18] extends ITI by introducing a non-linear multi-token probing and multi-token intervention method.
@@
  • Learnable Intervention for Truthfulness Optimization (LITO) [3] explores a sequence of model generations based on increasing levels of intervention magnitude and then selects the most accurate response.
Metrics. Following the standard benchmark in TruthfulQA [25, 24], we compare our method to baselines using the metrics described below.
@@
  • Two metrics for the multiple-choice task introduced in [25], namely MC1 and MC2. Given a question and some choices, select the only correct answer. The selection of the model is the answer choice to which it assigns the highest log-probability of completion following the question, independent of the other answer choices. The accuracy across all questions is denoted as MC1. Similarly, given a question and multiple true/false reference answers, the MC2 is the normalized total probability assigned to the set of true answers.
@@
  • For the generation task, we use two fine-tuned GPT-3.5-instruct models to classify whether an answer is true or false and informative or not. We report two metrics from [24]: truthful score True (%) and True*Info (%), a product of scalar truthful and informative score. We note that there are discrepancies between the results of ITI reproduced in our work and the original results reported in [24], as the original paper used GPT-3 based models to score these two metrics; however, at the time this paper is written, GPT-3 is no longer available on the OpenAI platform.
  • We report two additional metrics, Kullback-Leiber divergence (KL) of the model’s next-token prediction distribution post-versus-pre-intervention, and Cross-Entropy Loss (CE). These two metrics measure how much the generation distribution shifts after the intervention. Lower values are preferred since the intervention does not change the behavior of the original model dramatically and is unlikely to cause abnormal characters or non-natural sentences. The calculation of these metrics is elaborated in [24].
@@
Computing resources. We run all experiments on 4 NVIDIA RTX A5000 GPUs, an i9 14900K CPU, and 128GB RAM. The semidefinite programs (4) are solved using Mosek 10.1, with the average solving time for each instance being around 50 seconds.
Reproducibility. The repository is https://github.com/nguyenngocbaocmt02/OT-Intervention.
4.2. Numerical Results
4.2.1. Comparison between Finetuning-free Techniques
@@
We benchmark two fine-tuning-free baselines (ITI and FSP) along with our framework RADIANT on Llama-7B, Llama3-8B, and Llama2-chat-13B with the TruthfulQA dataset. The results are presented in Table 2. Across the three models, the combined method of FSP + RADIANT consistently achieved the highest scores in metrics such as True * Info and True, with 49% for Llama-7B, 44% for Llama3-8B, and 65% for Llama2-chat-13B. When running alone, our method, RADIANT, also demonstrated significant improvements, particularly in Llama2-chat-13B, where it achieved a True * Info score of 64% and a Truthful score of 74%. This suggests the efficiency of our framework compared with other baselines, including the current state-of-the-art ITI.
@@
Table 2. Quantitative results of different intervention methods on TruthfulQA dataset, across different Language Models. Parameters of RADIANT: $\alpha=2.5,\Gamma=15$.
Unintervened 21.15 22.16 25.58 40.54 2.13 0.00
SFT* 36.10 47.10 24.20 - 2.10 0.01
ITI 26.52 28.03 27.78 43.59 2.20 0.07
FSP 36.13 39.78 34.03 50.34 2.13 0.00
NL-ITI 29.06 38.04 32.97 45.69 2.19 0.07
LITO 39.08 41.22 29.22 47.64 2.19 0.07
RADIANT (ours) 40.36 44.48 30.91 46.13 2.19 0.07
FSP + ITI 40.63 45.16 35.50 52.48 2.20 0.07
FSP + NL-ITI 45.97 47.31 38.37 53.61 2.20 0.07
FSP + LITO 49.05 55.68 36.23 54.92 2.20 0.07
FSP + RADIANT (ours) 49.31 57.43 37.97 55.31 2.20 0.08
(a) Llama-7B
Unintervened 32.88 44.18 30.36 48.98 2.38 0.00
ITI 35.92 46.88 32.07 49.84 2.50 0.13
FSP 36.32 39.78 35.74 52.93 2.38 0.00
NL-ITI 35.98 45.72 33.02 51.37 2.50 0.13
LITO 37.53 48.20 34.96 52.54 2.48 0.11
RADIANT (ours) 37.78 50.82 33.82 52.98 2.48 0.08
FSP + ITI 40.63 45.16 35.50 52.98 2.48 0.14
FSP + NL-ITI 40.70 46.03 34.15 53.35 2.49 0.14
FSP + LITO 43.95 49.82 38.41 55.31 2.54 0.17
FSP + RADIANT (ours) 44.09 52.02 37.98 54.61 2.52 0.15
(b) Llama3-8B
Unintervened 51.87 59.86 35.38 53.32 2.31 0.00
ITI 57.02 63.04 37.46 55.59 2.32 0.17
FSP 55.97 58.63 40.76 57.84 2.31 0.00
NL-ITI 57.13 60.82 39.01 57.24 2.33 0.17
LITO 58.12 61.36 38.25 57.21 2.34 0.18
RADIANT (ours) 63.68 74.20 39.95 58.18 2.35 0.18
FSP + ITI 56.78 59.24 41.50 59.01 2.33 0.13
FSP + NL-ITI 59.62 61.77 42.15 57.87 2.34 0.15
FSP + LITO 60.74 63.21 41.28 58.46 2.36 0.17
FSP + RADIANT (ours) 64.68 67.75 42.52 59.99 2.38 0.18
(c) Llama2-chat-13B
4.2.2. Comparison between ITI, RADIANT, and Instruction Finetuning Methods.
@@
In this benchmark, we investigate whether implementing RADIANT on Alpaca and Vicuna, two instruction fine-tuning models from Llama-7B, can further enhance their performances. Results in Table 3 indicate that applying RADIANT significantly enhances both the baseline models, with Alpaca + RADIANT improved to 44.5% in True*Info score and 46% in Truthful score. Similarly, Vicuna + RADIANT achieved the highest scores of 55% in True*Info score and 63% in Truthful score, showcasing a marked increase compared to its baseline performance of 38% and 42.1%, respectively. In both cases, RADIANT outperformed ITI, demonstrating its effectiveness in enhancing the models’ accuracy and truthfulness.
4.3. Comparison between ITI, RADIANT, and Representation Intervention Finetuning Methods.
In this experiment, we apply RADIANT and ITI on Llama-7B, Llama3-8B, and Llama2-chat-13B models, which were previously fine-tuned by LOFiT, a representation intervention finetuning method. The experimental results in Table 4 show that RADIANT is better than ITI in improving correctness and informativeness across different Llama models. While ITI offers modest improvements in some instances, it generally lags behind RADIANT, especially in larger models. The KL divergence values suggest that RADIANT maintains a close distribution to the base model (LOFiT) while delivering substantial performance improvements.
4.3.1. Ablation study
We perform two ablation studies to demonstrate the effectiveness of our framework. In the first scenario, we select intervened heads using ITI, then compare our intervention approach versus ITI. In the second ablation study, the probing loss function is replaced by the widespread classification loss: the binary cross-entropy loss. Table 5 below reports the performance of the Llama-7B + TruthfulQA dataset. In the first scenario, switching the head selection between RADIANT and ITI improved performance when the RADIANT intervention was applied, reaching 37% in the True * Info score. The second scenario, which tested the impact of replacing the risk-aware loss function with cross-entropy loss, resulted in moderate improvements but still fell short compared to RADIANT’s risk-aware loss in Section 2 (30.36% vs 40.36% in True*Info). Overall, these findings highlight the effectiveness of our framework and suggest that both the choice of intervention and the loss function play crucial roles.
5. Conclusion
In this paper, we introduced RADIANT, a novel intervention framework for model editing consisting of two components: (i) a layerwise probe to detect undesirable content and (ii) headwise interventions to rectify the head activations upon undesirable-predicted outcome. Contrary to existing intervention methods, where the interventions can be scattered across different layers, our intervention is focused on a single layer of the network. This focus helps alleviate the distributional shifts of the activations in subsequent layers, which could reduce the performance of the detections and interventions therein. Moreover, our headwise intervention aims to minimize the perturbations to the activations while keeping a reasonable guarantee of the effectiveness of the intervention. This is further demonstrated in empirical results, where our method outperforms the state-of-the-art intervention method ITI [24] on various LMs.
Social Impact. Our paper focuses on improving the truthfulness of LMs, and the results aim to improve trustworthy artificial intelligence. Apart from language generation, our paper can also be implemented in other domains for activation editing. Nevertheless, it is important to acknowledge the potential misuse of our method. There exists a risk that adversarial actors could exploit our approach to transform truthful outputs into misleading or false information. This dual-use nature underscores the importance of ethical guidelines and safeguards in developing artificial intelligence. By promoting transparency and accountability in using our framework, we want to raise awareness of the risks while maximizing the benefits of improved truthfulness in language generation.
Table 3. Quantitative results of intervention methods on instruction-finetuned models Alpaca and Vicuna.
Alpaca 30.39 30.85 26.56 41.63 2.81 0.00
Alpaca + ITI 37.67 38.19 28.89 45.19 2.88 0.14
Alpaca + RADIANT (ours) 44.51 45.94 30.79 47.83 2.81 0.13
Vicuna 38.24 42.10 31.83 48.48 2.67 0.00
Vicuna + ITI 49.27 53.25 33.42 51.80 2.77 0.26
Vicuna + RADIANT (ours) 54.87 62.81 35.76 55.14 2.73 0.27
Table 4. Quantitative results of different intervention methods on TruthfulQA dataset, across different Language Models. We considered LOFiT as the base model for this experiment, so the KL of LOFiT is 0.
LOFiT 59.48 69.03 51.04 70.78 2.35 0.00
LOFiT + ITI 60.84 72.29 51.41 70.84 2.55 0.14
LOFiT + RADIANT (ours) 61.50 72.08 51.80 71.29 2.56 0.13
(a) Llama-7B
LOFiT 68.80 90.08 59.00 77.93 3,27 0.00
LOFiT + ITI 67.57 79.31 55.33 75.85 3.33 0.08
LOFiT + RADIANT (ours) 71.47 90.19 59.30 76.56 3.38 0.11
(b) Llama3-8B
LOFiT 66.35 81.89 57.04 76.17 2.52 0.00
LOFiT + ITI 66.00 78.09 55.08 75.25 2.73 0.21
LOFiT + RADIANT (ours) 69.63 83.86 57.45 75.47 2.73 0.20
(c) Llama2-chat-13B

### v2

@@ 3 Headwise Interventions with Probabilistic Guarantees
where $\Delta_{\ell h}(a_{\ell h})=G^{\star}_{\ell h}a_{\ell h}+g^{\star}_{\ell h}$ for $h=1,\ldots,H$. In other words, each new headwise activation $\hat{a}_{\ell h}$ is computed based on three terms: the original headwise activations $a_{\ell h}$, the headwise intervention $\Delta_{h}(a_{\ell h})$, and the indicator value identifying if head $h$ and layer $\ell$ is predicted desirable or undesirable.
4 Experiments
@@
In this section, we present empirical evidence for the effectiveness of our method RADIANT. We evaluate RADIANT on the TruthfulQA benchmark Lin et al. (2021), consisting of two tasks: the main task is the generation, and the secondary task is multiple choice. The generation task requires the model to generate an entire answer for each question using greedy autoregressive decoding. The accuracy and helpfulness of the answer are best assessed by humans. However, in almost all recent works in the field, including Li et al. (2024b) and Yin et al. (2024), this criterion is measured by an alternative large language model finetuned on the target dataset. The multiple-choice task contains candidate answers to each question, requiring the model to give probabilities. Higher probabilities for truthful answers yield higher scores.
4.1 Experimental Settings
@@
Datasets. We evaluate and compare our method with other baselines using the TruthfulQA benchmark Lin et al. (2021). Details about this dataset and how we preprocess the data can be found in Appendix A.1. In addition, we also show the generalization of our method by conducting a transferability experiment on two other out-of-distribution datasets, including NQOpen Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017). Due to space constraints, the results for the latter two datasets are relegated to Appendix A.5.
@@
Hyperparameters. There are two pivotal hyperparameters in the RADIANT framework, namely $\alpha$ in probe loss (2), and $\Gamma=\Phi^{-1}(1-\gamma)$ in the computation of the intervention map (4). The discussion about their impact on RADIANT and how to select them is in Appendix A.2.3.
Baselines. We benchmark against:
@@
  • Inference-time Intervention (ITI) Li et al. (2024b), the state-of-the-art method for finetuning-free intervention. The hyperparameters of the baseline follow their original paper and their GitHub repository. [^1] https://github.com/likenneth/honest_llama/tree/master
  • Few-shot prompting (FSP) introduced in Bai et al. (2022) showcases the effectiveness of 50-shot prompting in benchmark TruthfulQA.
  • Probe-Free Low Rank Activation Intervention (FLORAIN) Jiang et al. (2025) which trains a low rank component for intervening into one specified hidden representation.
@@
  • Instruction Fine-Tuning (IFT, Wang et al. 2022; Chung et al. 2024) is a popular fine-tuning approach to boost the truthfulness of language models. Two notable pretrained models in this direction, namely Alpaca-7B Taori et al. (2023) and Vicuna-7B Chiang et al. (2023), are adopted for comparison.
@@
  • Representation Intervention Fine-tuning (RIFT) methods aim to adjust language model activations for improved truthfulness. However, they add extra parameters and require extensive computational resources for fine-tuning. We consider LOFiT Yin et al. (2024) for comparison.
@@
  • Non-Linear Inference Time Intervention (NL-ITI) Hoscilowicz et al. (2024) extends ITI by introducing a non-linear multi-token probing and multi-token intervention method.
@@
  • Learnable Intervention for Truthfulness Optimization (LITO) Bayat et al. (2024) explores a sequence of model generations based on increasing levels of intervention magnitude and then selects the most accurate response.
Metrics. Following the standard benchmark in TruthfulQA Lin et al. (2021); Li et al. (2024b), we use the below metrics:
@@
  • For the multiple choice task, we use MC1 and MC2 metrics as defined in Lin et al. (2021). MC1 measures the model’s accuracy in selecting the correct answer from the given choices, where selection is based on the highest log-probability score assigned to each completion. MC2 is the normalized total probability assigned to the set of true answers.
@@
  • For the generation task, we use two fine-tuned GPT-3.5-instruct models to classify whether an answer is true or false and informative or not. We report two metrics from Li et al. (2024b): truthful score True (%) and True*Info (%), a product of scalar truthful and informative score. We note that there are discrepancies between the results of ITI reproduced in our work and the original results reported in Li et al. (2024b), as the original paper used GPT-3 based models to score these two metrics; however, at the time this paper is written, GPT-3 is no longer available on the OpenAI platform.
  • We assess how our method and baselines alter the original generation distribution using two extra metrics: the Kullback-Leibler (KL) divergence and Cross-Entropy (CE) loss of the model’s next-token prediction distribution before and after intervention. Due to limited space in the main paper, these metrics for our method and baselines are detailed in Appendix A.3.
@@
Computing resources. We run all experiments on 4 NVIDIA RTX A5000 GPUs, an i9 14900K CPU, and 128GB RAM. The semidefinite programs (4) are solved using Mosek 10.1; the average solving time for each instance is around 50 seconds.
4.2 Numerical Results
4.2.1 Comparison between Finetuning-free Techniques
@@
We benchmark two fine-tuning-free baselines (ITI and FSP) along with our framework RADIANT on Llama-7B, Llama3-8B, and Llama2-chat-13B with the TruthfulQA dataset. The results are presented in the first three big rows of Table 1. Across the three models, the combined method of FSP + RADIANT consistently achieved the highest scores in metrics such as True * Info and True, with 49% for Llama-7B, 44% for Llama3-8B, and 65% for Llama2-chat-13B. When running alone, our method, RADIANT, also demonstrated significant improvements, particularly in Llama2-chat-13B, where it achieved a True * Info score of 64% and a Truthful score of 74%. This suggests the efficiency of our framework compared with other baselines, including the current state-of-the-art ITI.
Model Methods True * Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Llama-7B Unintervened 21.15 22.16 25.58 40.54
ITI 26.52 28.03 27.78 43.59
FSP 36.13 39.78 34.03 50.34
NL-ITI 29.06 38.04 32.97 45.69
LITO 39.08 41.22 29.22 47.64
FLORAIN 31.46 34.72 31.76 47.43
RADIANT (ours) 40.36 44.48 30.91 46.13
FSP + ITI 40.63 45.16 35.50 52.48
FSP + NL-ITI 45.97 47.31 38.37 53.61
FSP + LITO 49.05 55.68 36.23 54.92
FSP + FLORAIN 45.31 49.23 36.45 54.27
FSP + RADIANT (ours) 49.31 57.43 37.97 55.31
Llama3-8B Unintervened 32.88 44.18 30.36 48.98
ITI 35.92 46.88 32.07 49.84
FSP 36.32 39.78 35.74 52.93
NL-ITI 35.98 45.72 33.02 51.37
LITO 37.53 48.20 34.96 52.54
FLORAIN 36.78 48.67 34.56 53.68
RADIANT (ours) 37.78 50.82 33.82 52.98
FSP + ITI 40.63 45.16 35.50 52.98
FSP + NL-ITI 40.70 46.03 34.15 53.35
FSP + LITO 43.95 49.82 38.41 55.31
FSP + FLORAIN 42.15 47.32 36.98 55.83
FSP + RADIANT (ours) 44.09 52.02 37.98 54.61
Llama2-chat-13B Unintervened 51.87 59.86 35.38 53.32
ITI 57.02 63.04 37.46 55.59
FSP 55.97 58.63 40.76 57.84
NL-ITI 57.13 60.82 39.01 57.24
LITO 58.12 61.36 38.25 57.21
FLORAIN 60.68 67.70 39.65 59.57
RADIANT (ours) 63.68 74.20 39.95 58.18
FSP + ITI 56.78 59.24 41.50 59.01
FSP + NL-ITI 59.62 61.77 42.15 57.87
FSP + LITO 60.74 63.21 41.28 58.46
FSP + FLORAIN 61.14 62.45 44.52 61.48
FSP + RADIANT (ours) 64.68 67.75 42.52 59.99
Alpaca Base 30.39 30.85 26.56 41.63
+ ITI 37.67 38.19 28.89 45.19
+ RADIANT (ours) 44.51 45.94 30.79 47.83
Vicuna Base 38.24 42.10 31.83 48.48
+ ITI 49.27 53.25 33.42 51.80
+ RADIANT (ours) 54.87 62.81 35.76 55.14
Llama variants + LOFiT LOFiT (7B) 59.48 69.03 51.04 70.78
+ ITI 60.84 72.29 51.41 70.84
+ RADIANT (ours) 61.50 72.08 51.80 71.29
LOFiT (8B) 68.80 90.08 59.00 77.93
+ ITI 67.57 79.31 55.33 75.85
+ RADIANT (ours) 71.47 90.19 59.30 76.56
LOFiT (Chat-13B) 66.35 81.89 57.04 76.17
+ ITI 66.00 78.09 55.08 75.25
+ RADIANT (ours) 69.63 83.86 57.45 75.47
@@
Table 1: Quantitative results of different intervention methods on TruthfulQA dataset, across different Language Models and fine-tuning approaches. Parameters of RADIANT: $\alpha=2.5,\Gamma=15$.
4.2.2 Comparison against ITI with Instruction Finetuning Methods.
@@
We investigate whether implementing RADIANT on Alpaca and Vicuna, two instruction fine-tuning models from Llama-7B, can further enhance their performances. Results in Table 1 (fourth and fifth big rows) indicate that applying RADIANT significantly enhances both the baseline models, with Alpaca + RADIANT improved to 44.5% in True*Info score and 46% in Truthful score. Similarly, Vicuna + RADIANT achieved the highest scores of 55% in True*Info score and 63% in Truthful score, showcasing a marked increase compared to its baseline performance of 38% and 42.1%, respectively. In both cases, RADIANT outperformed ITI, demonstrating its effectiveness in enhancing the models’ accuracy and truthfulness.
Methods True * Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Unintervened 21.15 22.16 25.58 40.54
ITI 26.52 28.03 27.78 43.59
1st scenario: Our linear probe + ITI intervention 26.88 28.00 29.00 44.00
1st scenario: ITI linear probe + our intervention 36.66 39.00 28.00 43.00
2nd scenario: Cross entropy loss 30.36 33.00 29.00 43.00
RADIANT 40.36 44.48 30.91 46.13